OpenStudy (anonymous):

if f(x)=1/9x-2, what is f^-1(x)

OpenStudy (anonymous):

You first switch f(x) with x in this equation \[x = \frac{ 1 }{ 9f(x) - 2 }\] Then you solve for f(x). \[9f(x) -2 = \frac{ 1 }{ x }\] \[9f(x) = \frac{ 1 }{ x } + 2\] \[f(x) = \frac{ 1 }{ 9x } + \frac{ 2 }{ 9 }\] Then you simply repalace f(x) with f^-1(x) It will look like this \[f^-1(x) = \frac{ 1 }{ 9x } + \frac{ 2 }{ 9 }\]